### v3

# Breaking the Modality Barrier: Universal Embedding Learning with Multimodal LLMs

Abstract.
The Contrastive Language-Image Pre-training (CLIP) framework has become a widely used approach for multimodal representation learning, particularly in image-text retrieval and clustering. However, its efficacy is constrained by three key limitations: (1) text token truncation, (2) isolated image-text encoding, and (3) deficient compositionality due to bag-of-words behavior. While recent Multimodal Large Language Models (MLLMs) have demonstrated significant advances in generalized vision-language understanding, their potential for learning transferable multimodal representations remains underexplored. In this work, we present UniME (Universal Multimodal Embedding), a novel two-stage framework that leverages MLLMs to learn discriminative representations for diverse downstream tasks. In the first stage, we perform textual discriminative knowledge distillation from a powerful LLM-based teacher model to enhance the embedding capability of the MLLM’s language component. In the second stage, we introduce hard negative enhanced instruction tuning to further advance discriminative representation learning. Specifically, we initially mitigate false negative contamination and then sample multiple hard negatives per instance within each batch, forcing the model to focus on challenging samples. This approach not only improves discriminative power but also enhances instruction-following ability in downstream tasks. We conduct extensive experiments on the MMEB benchmark and multiple retrieval tasks, including short&long caption retrieval and compositional retrieval. Results demonstrate that UniME achieves consistent performance improvement across all tasks, exhibiting superior discriminative and compositional capabilities.
Vision-Language Model, Multi-Modal, Compositional Understanding
* Equal Contribution
$\dagger$ Corresponding Author
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754845isbn: 979-8-4007-2035-2/2025/10
Figure 1. The UniME framework incorporates textual discriminative knowledge distillation and hard negative enhanced instruction tuning stages to learn discriminative representations for diverse downstream tasks. Our framework achieves state-of-the-art performance on both the MMEB benchmark and multiple retrieval tasks.
[…]
[¡short description¿]¡long description¿
1. Introduction
Modern AI applications increasingly rely on multimodal embeddings to process diverse data types, powering essential tasks like image-text retrieval (Baldrati et al., 2023; Tang et al., 2025), Retrieval Augmented Generation (RAG) (Jiang et al., 2023; Cong et al., 2023), and Visual Question Answering (VQA) (Gardères et al., 2020; Faysse et al., 2024; Chun et al., 2021). As a seminal model, CLIP (Radford et al., 2021) demonstrates notable text-image retrieval performance via cross-modal contrastive supervision using large-scale web-collected image-text pairs. However, despite its widespread use, CLIP presents notable limitations. Firstly, it restricts text token length to 77, hindering its ability to process detailed descriptions and limiting its utility in cross-modal retrieval tasks that require extensive contextual information (Zhang et al., 2024c; Cao et al., 2024; Huang et al., 2024). Moreover, CLIP employs a dual-encoder architecture that processes images and text separately, which compromises its effectiveness in complex tasks such as instruction-following multimodal retrieval (Jiang et al., 2025; Liu et al., 2024a; Wei et al., 2024). Additionally, CLIP exhibits limited advanced language understanding, struggles with compositionality, and tends to display bag-of-words behavior (Yuksekgonul et al., 2022; Tschannen et al., 2023).
The success of Large Language Models (LLMs) (Touvron et al., 2023a, b; Grattafiori et al., 2024; Bai et al., 2023; Yang et al., 2024) has motivated researchers to adapt LLMs to understand multimodal inputs. Multimodal Large Language Models (MLLMs) as a key component in the construction of general-purpose AI assistants have demonstrated remarkable progress (Liu et al., 2023, 2024b). For example, Qwen2-VL (Wang et al., 2024a) innovates beyond fixed-resolution visual processing, achieving robust performance across diverse image resolutions and aspect ratios. Similarly, LLaVA-OneVision (Li et al., 2024) introduces a unified modeling approach that enables effective task transfer across scenarios while maintaining architectural simplicity. While these MLLMs show impressive vision-language reasoning capabilities, these MLLMs are inherently constrained by their autoregressive next-token prediction objective, which limits their effectiveness in learning multimodal representations compared to contrastive methods such as CLIP (Jiang et al., 2024b, 2025).
Recent advances in LLM-based models have demonstrated substantial progress on the MTEB benchmark (Muennighoff et al., 2022). Inspired by these developments (Lee et al., 2024; BehnamGhader et al., 2024), researchers are now actively investigating MLLMs for unified multimodal representation learning. E5-V (Jiang et al., 2024b) proposes an unimodal contrastive learning approach that trains the language component of MLLM on sentence pairs to bridge cross-modal representation disparities. However, this method encounters two primary constraints: (1) constraints arising from the limited scale and diversity of training data (Ouali et al., 2024); (2) inherent challenges caused by the MLLM’s causal attention mechanism, which fundamentally restricts its ability to learn complex contextual representations (Man et al., 2024; Xie et al., 2024b, a). These factors collectively constrain the model’s full embedding potential. VLM2Vec (Jiang et al., 2025) introduces the Massive Multimodal Embedding Benchmark (MMEB), comprising 36 datasets across 4 meta-tasks, and develops a contrastive framework that converts state-of-the-art vision-language models into embedding models through MMEB training. Nevertheless, the existence of false negative samples in the batch significantly complicates the discrimination of hard negative pairs when using the standard InfoNCE loss.
To overcome these challenges, we present UniME (Universal Multimodal Embedding), a novel two-stage framework that empowers multimodal large language models (as shown in Figure 1) to learn universal representations for diverse downstream vision-language tasks. In the first textual discriminative knowledge distillation stage, we leverage a powerful LLM-based teacher model to enhance the embedding capabilities of MLLM’s language component. In the second stage of hard negative enhanced instruction tuning, we first eliminate false negative contamination, then implement a hard negative sampling strategy that selects multiple challenging negatives per instance within each batch. This approach forces the model to focus on challenging negative samples, thereby learning more discriminative multimodal representations while also improving instruction-following ability in downstream tasks. We evaluate our approach comprehensively on the MMEB benchmark and multiple retrieval tasks, including both short&long caption retrieval and compositional retrieval. Experimental results demonstrate that UniME achieves significant performance improvement across all tasks, exhibiting both robust discriminative power and superior compositional understanding. The main contributions of this paper are summarized as follows:
  • We present UniME (Universal Multimodal Embedding), a novel two-stage framework that empowers Multimodal Large Language Models (MLLMs) to learn universal representations for diverse downstream tasks.
  • We propose textual discriminative knowledge distillation, leveraging a powerful LLM-based teacher model to enhance the embedding capability of the MLLM’s language component.
  • We introduce hard negative enhanced instruction tuning to further advance discriminative representation learning through false negative filtering and hard negative sampling.
  • We conduct extensive experiments on the MMEB benchmark and multiple retrieval tasks, including both short&long caption retrieval and compositional retrieval. Results show that UniME demonstrates robust discriminative and compositional capabilities, achieving notable performance improvements across all evaluated tasks.
2. Related work
Figure 2. The framework of the Textual Discriminative Knowledge Distillation stage. We leverage the state-of-the-art LLM-based embedding model to enhance the discriminative capabilities of the MLLM’s language component.
[…]
[¡short description¿]¡long description¿
2.1. Multimodal Large Language Models
Multimodal Large Language Models (MLLMs) extend LLMs to process and integrate cross-modal information (Bin et al., 2024; Wang et al., 2024b, b; Jiang et al., 2024a). A seminal development in this field is LLaVA (Liu et al., 2023), which encodes images into token representations using vision encoders like CLIP (Radford et al., 2021; Sun et al., 2023a; Yang et al., 2023; Gu et al., 2024; Yang et al., 2025; Hu et al., 2025) and projects them into the LLM’s token space. Following this breakthrough, numerous MLLM variants (Peng et al., 2023; Lin et al., 2024; Zhang et al., 2024a; Zhu et al., 2023) have demonstrated remarkable performance in multimodal comprehension and reasoning tasks. CogVLM (Wang et al., 2023) incorporates a trainable visual expert module within the attention and feed-forward network layers of the language model, yielding substantial performance enhancements across 17 canonical cross-modal benchmarks. LLaVA-1.6 (Liu et al., 2024c) introduces the “AnyRes” technique to process variable high-resolution images, significantly improving fine-grained visual understanding. Phi3.5-Vision (Abdin et al., 2024), a 4.2-billion parameter model evolved from phi-3.5-mini, exhibits strong reasoning capabilities for both single and multi-image text prompts. While these advancements, the autoregressive next-token prediction objective of MLLM inherently constrains its capacity to learn efficient multimodal representations.
$2.$ LLMs for Representation Learning
As large language models increasingly exhibit remarkable proficiency in natural language processing, recent research has pivoted towards harnessing decoder-only architectures for effective representation learning (Ma et al., 2024; Lee et al., 2024; BehnamGhader et al., 2024; Shin et al., 2025; Didolkar et al., 2025). Previous work has adapted the prompt-based representation method for autoregressive models, enabling Large Language Models (LLMs) to perform in-context learning and scale to various model sizes. LLM2Vec (BehnamGhader et al., 2024) transforming pre-trained decoder-only LLMs into versatile text encoders by incorporating three principal advancements: bidirectional attention mechanisms, masked next-token prediction, and unsupervised contrastive alignment. Concurrently, NV-Embed (Lee et al., 2024) introduces a latent attention layer and eliminates the causal attention mask during contrastive training, substantially enhancing the efficiency of embeddings generated from decoder-only LLMs. While these approaches show promising embedding performance, their exclusive focus on text-only inputs fails to meet the growing demands of multimodal applications.
2.3. Multimodal Representation Learning
CLIP (Radford et al., 2021) demonstrates superior image-text retrieval capabilities through large-scale cross-modal contrastive learning, but it suffers from three inherent constraints: (1) 77-token text truncation limits fine-grained semantic alignment (Zhang et al., 2024c; Cao et al., 2024; Huang et al., 2024); (2) Disjoint dual-encoder architecture impedes cross-modal fusion for instruction-sensitive tasks (Jiang et al., 2025; Liu et al., 2024a; Wei et al., 2024); (3) Primitive language modeling induces bag-of-words representations (Yuksekgonul et al., 2022; Tschannen et al., 2023). Recent research addresses these limitations through two complementary approaches. Firstly, MagicLens (Zhang et al., 2024b) employs lightweight dual encoders to enable relation-aware image retrieval guided by textual instructions. Secondly, MLLM-based approaches emerge for robust representation learning. E5-V (Jiang et al., 2024b) proposes a single-modality training approach that significantly outperforms traditional multimodal training on image-text pairs while reducing training costs. VLM2Vec (Jiang et al., 2025) presents a contrastive training framework that can handle any combination of images and text, as well as high-resolution images and long text inputs. Despite these improvements, current methods still face challenges in effectively discriminating hard negative samples during retrieval.
3. Method
Figure 3. The framework of the Hard Negative Enhanced Instruction Tuning stage. We further improve the discriminative capabilities of the MLLM through false negative filtering and hard negative sampling.
[…]
[¡short description¿]¡long description¿
This section first establishes the preliminary definitions, including task formulation and feature extraction (Section 3.1). Then we start to introduce our proposed novel two-stage framework UniME. We elaborate on the first textual discriminative knowledge distillation stage in Section 3.2. After that, we present the second hard negative enhanced instruction tuning stage in Section $3.$ Finally, we introduce the training recipe in Section 3.4.
3.1. Preliminary
3.1.1. Task Definition.
To address the limitations of dual-tower encoder structures in acquiring unified multimodal representations (Radford et al., 2021), we employ Multimodal Large Language Models (MLLMs) with robust multimodal understanding to learn universal multimodal embeddings. Specifically, we feed both the query and candidate data into the MLLM using customized prompts to extract their respective embeddings. We then calculate the similarity ($\Theta$) between the query ($\boldsymbol{q}$) and candidates ($\boldsymbol{c}$), followed by ranking and selecting the most relevant pairs ($\boldsymbol{P}$). This procedure is formalized as follows:
(1) $\boldsymbol{P}=\text{Rank}(\Theta(\phi(\boldsymbol{q}),\phi(\boldsymbol{c}))),$
where $\phi$ denotes the MLLM employed to extract respective embeddings. Notably, both the query and candidate may be either unimodal (text or image only) or multimodal (interleaved image-text).
3.1.2. Feature Extraction by MLLM.
[…]
3.2. Textual Discriminative Knowledge Distillation
3.2.1. Training.
[…]
(2) $\mathcal{L}_{KL}=\sum_{i=1}^{n}\mathrm{KL}\left(\frac{\exp\left(e_{s_{i}}^{\top}e_{s_{i}}/\tau\right)}{\sum\limits_{j=1}^{n}\exp\left(e_{s_{j}}^{\top}e_{s_{i}}/\tau\right)}\,\Bigg{\|}\,\frac{\exp\left(e_{t_{i}}^{\top}e_{t_{i}}/\tau\right)}{\sum\limits_{j=1}^{n}\exp\left(e_{t_{j}}^{\top}e_{t_{i}}/\tau\right)}\right),$
where $\tau$ is the temperature hyper-parameter used to soften distribution representation. By distilling the relationships between different samples within a batch, our method demonstrates enhanced efficiency compared to direct contrastive learning under identical data and training conditions and achieves significant performance improvements in downstream tasks.
3.2.2. Inference.
During the training phase, our approach exclusively utilizes text-only inputs and optimizes solely the language model component within the multimodal language model architecture, while maintaining other parameters frozen. For inference, we restore the original vision encoder and projection layer to enable multimodal processing. For unimodal inputs (text or image), we use modality-specific standardized prompts. For interleaved image-text inputs, we process each modality independently with its corresponding prompt and aggregate the embeddings through element-wise summation to produce the final multimodal representation.
Table 1. Results on the MMEB benchmark. IND represents the in-distribution dataset, and OOD represents the out-of-distribution dataset. The reported scores are the average Precision@1 over the corresponding datasets. The best results are marked in bold. : UniME with textual discrimination distillation only. : UniME with both textual discrimination distillation and hard negative enhanced instruction tuning.
Models #Parameters Per Meta-Task Score Average Score
Classification VQA Retrieval Grounding IND OOD Overall
# of Datasets $\rightarrow$ 10 10 12 4 20 16 36
Zero-shot on MMEB
CLIP(ViT-L) (Jiang et al., 2025) 0.4B 42.8 9.1 53.0 51.8 37.1 38.7 39.2
OpenCLIP(ViT-L) (Radford et al., 2021) 0.4B 41.5 6.9 44.6 53.5 32.8 36.0 36.6
Magiclens(ViT-L) (Zhang et al., 2024b) 0.4B 38.8 8.3 35.4 26.0 31.0 23.7 27.1
SigLIP(So/14) (Zhai et al., 2023) 0.9B 40.3 8.4 31.6 59.5 32.3 38.0 35.0
BLIP2(ViT-L) (Li et al., 2023) 1.2B 27.0 4.2 33.9 47.0 25.3 25.1 28.0
CLIP(ViT-BigG/14) (Cherti et al., 2022) 2.5B 52.3 14.0 50.5 60.3 38.9 45.8 44.3
EVA-CLIP (Sun et al., 2023b) 8B 56.0 10.4 49.2 58.9 38.1 45.6 43.7
E5-V(Phi3.5-V) (Jiang et al., 2024b) 4.2B 39.1 9.6 38.0 57.6 33.1 31.9 36.1
E5-V(LLaVA-1.6) (Jiang et al., 2024b) 7B 39.7 10.8 39.4 60.2 34.2 33.4 37.5
UniME(Phi3.5-V) 4.2B 42.5(+3.4) 18.3(+8.7) 40.5(+2.5) 59.9(+2.3) 36.0(+2.9) 38.3(+6.4) 40.3(+4.2)
UniME(LLaVA-1.6) 7B 43.0(+3.3) 17.7(+6.9) 42.5(+3.1) 63.2(+3.0) 37.6(+3.4) 38.6(+5.2) 41.6(+4.1)
Fine-tuning on MMEB
CLIP(ViT-L) (Jiang et al., 2025) 0.4B 55.2 19.7 53.2 62.2 47.6 42.8 47.6
VLM2Vec(Phi3.5-V) (Jiang et al., 2025) 4.2B 54.8 54.9 62.3 79.5 66.5 52.0 62.9
VLM2Vec(LLaVA-1.6) (Jiang et al., 2025) 7B 56.8 50.4 63.3 82.6 64.9 53.9 63.3
UniME(Phi3.5-V) 4.2B 54.8(+0.0) 55.9(+1.0) 64.5(+2.2) 81.8(+2.3) 68.2(+1.7) 52.7(+0.7) 64.2(+1.3)
UniME(LLaVA-1.6) 7B 60.6(+3.8) 52.9(+2.5) 67.9(+4.6) 85.1(+2.5) 68.4(+3.5) 57.9(+4.0) 66.6(+3.3)
UniME(LLaVA-OneVision) 7B 66.8 66.6 70.6 90.9 74.6 65.8 70.7
$3.$ Hard Negative Enhanced Instruction Tuning
After textual discriminative knowledge distillation, UniME develops preliminary discriminative capabilities but exhibits limited visual sensitivity. This insensitivity results in deviations in image-text alignment and limits the discriminative performance, despite the MLLM’s extensive pretraining on large-scale datasets. Moreover, the generic instruction prompts used in the first stage hinder UniME’s effectiveness in complex retrieval tasks. To address these limitations, we introduce an additional hard negative enhanced instruction tuning stage (as shown in Figure 3), which aims to: (1) further enhance discriminative capabilities, (2) improve cross-modal alignment, and (3) strengthen instruction-following ability for downstream tasks.
3.3.1. False Negative Contamination.
The presence of false negatives in training batches hinders effective hard negative differentiation under the standard InfoNCE loss. As evidenced in Table 8, false negatives frequently appear as candidate samples. To mitigate this, we introduce a filtering mechanism based on the similarity threshold between the query and positive samples, defined as: $\alpha=cos(e_{q},e_{c}^{+})+\beta$, where $\beta$ is a hyper-parameter used to control the threshold margin. During training, we exclude all negative samples whose similarity to the query exceeds $\alpha$, effectively eliminating false negatives while preserving challenging hard negatives.
3.3.2. Hard Negative Sampling
Hard negative samples, distinct in label from the positive sample but closely embedded, are poised to offer the greatest utility through the provision of substantial gradient information in the context of contrastive learning. In contrast, easy negatives yield negligible gradients and contribute minimally to the learning process. Drawing inspiration from prior research (Karpukhin et al., 2020; Robinson et al., 2020; Kalantidis et al., 2020), we propose a hard negative sampling strategy to optimize both training efficiency and discriminative performance. The textual knowledge distillation stage equips UniME with preliminary discriminative ability to autonomously identify hard negatives for each query. Building on this capability, we sample $k$ corresponding hard negative $e^{-}_{c}$ within each training batch as follows:
(3) $e^{-}_{c}=\text{Rank}_{k}(cos(e_{q},e_{c})),\text{where}\ e_{c}\notin\left\{e^{+}_{c},e^{*}_{c}\right\},$
where $e^{*}_{c}$ and $e^{+}_{c}$ denote filtered false negative candidates and positive candidates respectively, $e_{q}$ is the query embedding, and $e_{c}$ represents all candidate embeddings. The function $cos(\cdot)$ computes pairwise similarity scores, and $\text{Rank}_{k}$ selects the top-$k$ highest-scoring candidates as hard negatives. To maintain batch consistency when fewer than $k$ hard negatives are obtained (which occurs rarely), we duplicate existing hard negatives to preserve the fixed size $k$.
3.3.3. Training Objective.
After obtaining the embedding of the queries ($e_{q}$), positive candidates ($e_{c}^{+}$) and hard negative candidates ($e_{c}^{-}$), we utilize the Noise Contrastive Estimation (InfoNCE) loss (Oord et al., 2018) over the in-batch sampled hard negatives as follows:
(4) $\mathcal{L}=-log\frac{exp(cos(e_{q},e_{c}^{+})/\tau)}{exp(cos(e_{q},e_{c}^{+})/\tau)+\sum_{i=1}^{k}exp(cos(e_{q},e_{c_{i}^{-}})/\tau)},$
where $k$ denotes the set of all hard negatives, and $\tau$ is a temperature hyper-parameter.
Table 2. Results of zero-shot text-image retrieval on short caption datasets (Flickr30K and MS-COCO), long caption datasets (ShareGPT4V and Urban1K) and compositional benchmark (SugarCrepe). The reported scores are the average Recall@1 over the corresponding datasets. The best results are marked in bold. : UniME with textual discrimination distillation only. : UniME with both textual discrimination distillation and hard negative enhanced instruction tuning.
Models #Parameters Short Caption Retrieval Long Caption Retrieval Compositional Retrieval
Flickr30K COCO ShareGPT4V Urban1K SugarCrepe
$q^{i}\rightarrow c^{t}$ $q^{t}\rightarrow c^{i}$ $q^{i}\rightarrow c^{t}$ $q^{t}\rightarrow c^{i}$ $q^{i}\rightarrow c^{t}$ $q^{t}\rightarrow c^{i}$ $q^{i}\rightarrow c^{t}$ $q^{t}\rightarrow c^{i}$ Replace Swap Add
OpenCLIP(ViT-L) (Radford et al., 2021) 0.4B 67.3 87.2 37.0 58.1 81.8 84.0 47.0 47.0 79.5 62.7 74.9
CLIP(ViT-BigG/14) (Cherti et al., 2022) 2.5B 79.5 92.9 51.3 67.3 90.1 93.6 77.8 80.7 86.5 68.9 88.4
EVA-CLIP (Sun et al., 2023b) 8B 80.3 94.5 52.0 70.1 93.1 91.2 80.4 77.8 85.9 70.3 86.7
E5-V(Phi3.5-V) (Jiang et al., 2024b) 4.2B 72.2 79.6 44.7 53.4 86.0 88.5 83.8 83.6 88.2 66.6 75.3
E5-V(LLaVA-1.6) (Jiang et al., 2024b) 7B 77.3 85.7 49.1 57.6 85.1 82.1 88.9 83.2 86.3 68.7 66.9
UniME(Phi3.5-V) 4.2B 72.0(-0.2) 80.6(+1.0) 44.9(+0.2) 57.2(+0.8) 86.8(+3.8) 92.3(+1.3) 85.1(+2.3) 86.9(+3.3) 90.2(+2.0) 67.6(+1.0) 91.2(+15.9)
UniME(LLaVA-1.6) 7B 77.2(-0.1) 84.6(-1.1) 51.0(+1.9) 56.4(-1.2) 89.8(+4.7) 86.9(+4.8) 91.3(+2.4) 82.4(-0.8) 89.5(+3.2) 64.8(-3.9) 94.2(+27.3)
VLM2Vec(Phi3.5-V) (Jiang et al., 2025) 4.2B 68.7 83.0 43.7 59.8 90.1 92.0 87.9 86.8 86.2 66.7 84.2
VLM2Vec(LLaVA-1.6) (Jiang et al., 2025) 7B 76.0 90.6 46.8 66.6 85.8 90.7 84.7 90.8 85.8 66.3 86.5
UniME (Phi3.5-V) 4.2B 77.0(+11.3) 88.2(+5.2) 49.8(+6.1) 66.8(+7.0) 92.1(+2.0) 96.4(+4.4) 92.7(+4.8) 95.1(+8.3) 90.1(+3.9) 70.9(+4.2) 93.3(+9.1)
UniME (LLaVA-1.6) 7B 81.9(+5.9) 93.4(+2.8) 53.7(+6.1) 70.1(+3.5) 93.9(+8.1) 97.2(+6.5) 95.2(+10.5) 95.9(+5.1) 89.0(+3.2) 71.5(+5.2) 94.4(+7.9)
UniME (LLaVA-OV) 7B 83.3 94.4 54.8 74.0 93.9 89.3 94.3 95.5 80.5 65.5 82.2
3.4. Training Recipe
Stage1: Textual Discriminative Knowledge Distillation. We employ QLoRA (Quantized Low-Rank Adaptation) (Dettmers et al., 2023) for parameter-efficient fine-tuning of the large language model component. This stage exclusively utilizes text-only inputs and introduces minimal trainable parameters, typically less than 5% of the total. The complete training procedures for Phi3.5-V and LLaVA-1.6 require approximately 1 hour and 2 hours, respectively.
Stage2: Hard Negative Enhanced Instruction Tuning. To overcome GPU memory limitations in large-batch MLLM training, we employ a dual strategy: (1) Following VLM2Vec (Jiang et al., 2025), we implement GradCache (Gao et al., 2021b), a gradient caching technique that decouples backpropagation between contrastive loss computation and encoder updates; (2) we employ QLoRA (Dettmers et al., 2023) for parameter-efficient fine-tuning of all parameters within the MLLM. This combined approach facilitates effective training while ensuring manageable memory consumption.
4. Experiments
4.1. Implementation
[…]
4.2. Datasets and Metrics
4.2.1. Training.
Following E5-V (Jiang et al., 2024b), we utilize the Natural Language Inference (NLI) dataset (Gao et al., 2021a) which contains around 273k sentence pairs for the first textual discriminative knowledge distillation stage. For the hard negative enhanced instruction tuning stage, similar to the VLM2Vec (Jiang et al., 2025), we employ 20 in-distribution datasets from the MMEB benchmark, which cover four core multimodal tasks: classification, visual question answering, multimodal retrieval, and visual grounding. This comprehensive training corpus, incorporating both unimodal and multimodal input data, totals 662k carefully curated training pairs, ensuring robust model adaptation across diverse multimodal tasks.
4.2.2. Evaluation.
In this study, we evaluate UniME across both in-distribution (20 test sets) and out-of-distribution (16 test sets) benchmarks from MMEB (Jiang et al., 2025) to assess its multimodal embedding capabilities across diverse retrieval tasks. Following standard evaluation protocols (Liu et al., 2024a; Jiang et al., 2025), we report Precision, which quantifies the proportion of correct matches among the top-ranked candidates for each dataset. To further examine the unimodal embedding performance of UniME, we conduct experiments on multiple cross-modal retrieval tasks, including short-caption image-text retrieval (Flickr30K (Plummer et al., 2015) and COCO2014 (Lin et al., 2014)), long-caption image-text retrieval (ShareGPT4V (Chen et al., 2024) and Urban1K (Zhang et al., 2024c)), and compositional retrieval (SugarCrepe (Hsieh et al., 2023)). Consistent with the MMEB benchmark, Precision serves as the primary evaluation metric across all datasets.
4.3. Main Results.
4.3.1. Multi-Modal Retrieval
In Table 1, we present the performance of our proposed UniME against existing baseline models. Under identical training data and configuration settings, our proposed UniME consistently achieves significant performance improvements over E5-V across different foundation models. Specifically, UniME exhibits an average performance improvement of 4.2% over E5-V using the Phi3.5-V model. With LLaVA-1.6 as the base model, UniME further achieves an average enhancement of 4.1%. These significant performance improvements are primarily due to our proposed textual discrimination knowledge distillation method, which more efficiently enhances the discriminative capability of embeddings from a powerful teacher model. As depicted in Figure 4, we randomly select 50 samples from COCO and visualize the cross-modal cosine similarity matrix. The diagonal clarity of the UniME matrix is significantly enhanced compared to that of E5-V, indicating that UniME learns representations with superior distinctiveness. After the hard negative enhanced instruction tuning stage, beneficial from the hard negatives, the discriminative capability of the embeddings from UniME is further improved. Compared with VLM2Vec (Jiang et al., 2025), our model achieves 1.3% and 10.3% performance improvement when using Phi3.5-V and LLaVA-1.6 as the base model.
Figure 4. The discrimination comparison between E5-V and UniME. represents the UniME model only training on the first textual discrimination knowledge distillation stage.
[…]
[¡short description¿]¡long description¿
4.3.2. Short&Long Caption Cross-Modal Retrieval.
We evaluate our approach on zero-shot cross-modal retrieval tasks. Firstly, we conduct experiments on the short-caption datasets Flickr30K and MS-COCO. After the textual discrimination knowledge distillation stage, UniME achieves comparable retrieval performance to E5-V. Subsequent hard negative enhanced instruction tuning further boosts UniME’s performance, yielding a significant improvement of 5.2%–11.3% over VLM2Vec. For long-caption retrieval tasks on the ShareGPT4V and Urban1K datasets, UniME demonstrates superior performance across all metrics. Specifically, following the textual discriminative knowledge distillation stage, UniME exhibits a performance improvement of 1.3%-3.8% based on the Phi3.5-V model. Subsequent enhancement through hard negative enhanced instruction tuning leads to additional gains, with UniME outperforming VLM2Vec by 2.0%-8.3%. It is noteworthy that, compared to EVA-CLIP(8B), UniME achieves performance improvements of 14.8% and 18.1% in long-caption retrieval on the Urban1K dataset. This significant enhancement primarily stems from the limitation of EVA-CLIP(8B), which is constrained by a 77-token length restriction, thereby inhibiting its ability to fully convey the complete semantic information of long captions.
4.3.3. Compositional Cross-Modal Retrieval.
We evaluate the capacity of our UniME model to discriminate hard negative samples using the compositional benchmark SugarCrepe. As shown in Table 2, UniME consistently delivers superior performance across all evaluated metrics. Following the textual discriminative knowledge distillation phase, the Phi3.5-V-based UniME outperforms E5-V by 2.0% in relation replacement, 1.0% in object swapping, and 15.9% in attribute addition tasks. After the second hard negative enhanced instruction tuning stage, beneficial from the hard negative, the compositional understanding capabilities of UniME are further enhanced and it achieves 3.9%, 4.2%, and 9.1% performance improvement compared with VLM2Vec. Additionally, UniME exhibits improvements of 4.2%, 0.6%, and 6.6% compared to EVA-CLIP(8B), underscoring its robust capability to discriminate against hard negative samples.
5. Analysis
5.1. Analysis of the Hard Negative
In Figure 5, we visualize the training loss and pre-clip gradient norms for three negative types: easy negatives (least similar in batch), random negatives (randomly sampled in batch), and hard negatives (most similar in batch after removing positives and false negatives). Since easy negatives are easily distinguishable, the model struggles to enhance its discriminative power through learning from such data, consequently leading to a rapid convergence of the training loss to nearly zero. Using random negatives, the training loss converges more slowly compared to easy negatives, but it eventually nears zero. In contrast, hard negatives pose considerable challenges, sustaining elevated training losses. Correspondingly, gradient norms for easy negatives are minimal, whereas those for hard negatives are substantially higher, differing by orders of magnitude.
Figure 5. The comparison of training loss and pre-clip gradient norms for hard negatives, easy negatives, and random sample negatives.
[…]
[¡short description¿]¡long description¿
Figure 6. Visualization of the top-k next predicted tokens before and after different training stages based on Phi3.5-V. : UniME with textual discrimination distillation only. : UniME with both textual discrimination distillation and hard negative enhanced instruction tuning.
[…]
[¡short description¿]¡long description¿
Table 3. Ablation study of different training stages. We report the mean scores on the MMEB benchmark, short and long cross-modal retrieval, as well as compositional cross-modal retrieval.
Stage1 Stage2 MMEB $\text{R}_{\text{Short}}$ $\text{R}_{\text{Long}}$ $\text{R}_{\text{Compos}}$
25.3 44.2 62.9 63.1
40.3 63.7 87.8 83.0
63.8 61.5 84.2 77.1
64.2 70.4 94.1 84.8
5.2. Ablation on Training Stages
We conduct an ablation study based on Phi3.5-V to evaluate different training stages. As shown in Table 3, the initial embeddings from Phi3.5-V exhibit weak discriminative properties, leading to suboptimal task performance. After the initial stage of textual discriminative knowledge distillation, the model registers performance improvements of 15%, 19.5%, 24.9%, and 19.9% on the MMEB benchmark, short and long caption cross-modal retrieval, and compositional retrieval tasks, respectively. Focusing solely on the second stage, which involves hard negative enhanced instruction tuning, results in performance gains of 38.5%, 17.3%, 21.3%, and 14.0% in the same tasks. Notably, the enhancement in MMEB benchmark performance after the second stage markedly exceeds that of the first, primarily due to improved model capabilities in following complex instructions. By integrating both training stages, our UniME model achieves optimal performance across all evaluated downstream tasks.
5.3. Ablation on the Threshold $\beta$
The threshold $\beta$ directly influences the percentage of the filtered false negatives. We conduct an ablation study on 20% of the entire dataset to identify the optimal value of $\beta$. As illustrated in Table 8, setting $\beta$ to -0.1 results in filtering false negatives for 81.7% of samples. However, the inclusion of some hard negative samples in the filtered set results in poor model performance. Increasing $\beta$ from -0.1 to 0.1 reduces the proportion of samples with filtered false negatives from 81.7% to 22.9%, thereby significantly improving performance. Further increasing $\beta$ to 0.3 filters false negatives in only 13.1% of samples, resulting in a slight performance decline due to persistent false negatives.
Table 4. Ablation study of the false negative filtering threshold $\beta$. FalseNeg(%): proportion of samples which filtered false negatives.
Model $\beta$ FalseNeg(%) Average Score
IND OOD Overall
Phi3.5-V -0.1 81.7% 61.0 43.4 55.3
0.0 53.2% 66.1 49.0 61.1
0.1 22.9% 68.2 52.7 64.2
0.2 18.2% 68.2 51.9 63.7
0.3 13.1% 68.3 52.1 63.9
5.4. Ablation on the Number of Hard Negatives
The value of $k$ directly determines the number of hard samples sampled for each instance in a batch. In Table 5, we illustrates the impact of ($k$) on the performance of UniME based on Phi-3.5V. When we set kk=8, compared to kk=4, the higher diversity of hard negative samples leads to superior performance. Further increasing the value of $k$ introduces easy negatives, causing the model to lose focus on learning hard negatives, gradually reducing performance.
$5.$ Visualization of the Output Distribution
[…]
Table 5. Ablation study of the number $k$ of the hard negatives. HardNeg(%): proportion of hard negative samples within a batch.
Model Top-$k$ HardNeg(%) Average Score
IND OOD Overall
Phi3.5-V 4 0.4% 67.8 52.4 63.8
8 0.8% 68.2 52.7 64.2
16 1.6% 68.1 51.6 63.5
32 3.2% 68.4 51.8 63.6
64 6.4% 68.1 51.1 63.2
6. Conclusion
In this paper, we introduce UniME (Universal Multimodal Embedding), a novel two-stage framework that enables large multimodal language models with the capacity to learn discriminative representations applicable to a variety of downstream tasks. In the first textual discriminative knowledge distillation stage, we leverage a powerful LLM-based teacher model to enhance the embedding capability of the MLLM’s language component. In the second hard negative enhanced instruction tuning stage, we initially mitigate false negative contamination and then sample multiple hard negatives per instance within each batch, forcing the model to focus on challenging samples. This approach not only improves discriminative power but also enhances instruction-following ability in downstream tasks. We conduct extensive experiments on the MMEB benchmark and multiple retrieval tasks, including short&long caption retrieval and compositional retrieval. Results demonstrate that UniME achieves consistent performance improvement across all tasks, exhibiting superior discriminative and compositional capabilities. We hope that our work provides insights into multimodal representation learning.
Appendix A Detailed Experiment Settings
A.1. Training Details
We provide the training configurations of UniME in Table 6.
[…]
Stage2: Hard Negative Enhanced Instruction Tuning. Upon integrating the backbone with stage 1 LoRA outputs, we further advance UniME on image-text pairs derived from the MMEB training set. Under identical hardware conditions, this phase requires 26 hours for Phi-3.5V and 37 hours for LLaVA-1.6.
A.2. Retrieval Task Evaluation Benchmarks
We evaluate UniME on diverse retrieval benchmarks, including short-caption, long-caption, and compositional image-text tasks (as shown in Table 7). For each benchmark, we adopt its standard evaluation protocol. In retrieval tasks, we primarily report Recall@1 as the evaluation metric.
Table 7. Summary of the evaluation benchmarks. # Queries represents the number of test queries, and # Candidates denotes the number of test candidates per query.
Benchmark Zero-shot #Queries #Candidates
Flickr30K (Plummer et al., 2015) 1K 5K
COCO (Lin et al., 2014) 5K 25K
ShareGPT4V (Chen et al., 2024) 1K 1K
Urban1K (Zhang et al., 2024c) 1K 1K
SugarCrepe (Hsieh et al., 2023) 7.5K 2
Appendix B External Results
Table 8. Performance comparison of different LoRA ranks under two training stages (IND: In-Domain, OOD: Out-Of-Domain)
Stage LoRA Average Score
IND OOD Overall
Stage1 4 36.0 37.9 40.0
8 35.8 37.8 39.9
16 35.9 38.0 40.0
32 36.0 38.3 40.3
64 35.6 37.8 39.9
Stage2 4 67.4 51.0 62.6
8 68.2 51.0 63.8
16 68.2 52.7 64.2
32 67.6 51.8 63.2
64 67.0 51.5 62.7
Figure 7. Visualization of the top-k next predicted tokens before and after different training stages based on Phi3.5-V. : UniME with textual discrimination distillation only. : UniME with both textual discrimination distillation and hard negative enhanced instruction tuning.
[…]
[¡short description¿]¡long description¿
B.1. Ablation on the LoRA Rank
Due to computational limitations and the proven efficacy of LoRA (Dettmers et al., 2023), we utilize QLoRA to fine-tune the backbone of UniME. As detailed in Table 8, we assess the impact of various LoRA ranks on the MMEB benchmarks (Jiang et al., 2025) utilizing the Phi-3.5V backbone. Our findings indicate that a LoRA rank of 32 delivers optimal performance in Stage 1, whereas a rank of 16 is most effective in Stage 2.
B.2. Specific results on the MMEB
We present comparative results of eight models on the MMEB benchmark in Table 9. The performance metrics for CLIP, SigLIP, BLIP-2, and MagicLens are sourced directly from VLM2Vec (Jiang et al., 2025). Conversely, results for EVA-CLIP(8B), E5-V, VLM2Vec, and UniME are obtained through reproduction in our experiments. For E5-V, VLM2Vec, and UniME, we only report the metrics for their best-performing variants, all of which employ the LLaVA-1.6 (Liu et al., 2024c) backbone.
Figure 8. Schematic illustration of Negative data.
[…]
[¡short description¿]¡long description¿
Appendix C Further Analysis
Table 9. The detailed results of the baselines and our UniMEon MMEB, which includes 20 in-distribution datasets and 16 out-of-distribution datasets. The out-of-distribution datasets are highlighted with a yellow background in the table. We only introduce the best version of UniME, E5-V and VLM2Vec in the table, which uses LLaVA-1.6 as the backbone.
CLIP SigLIP BLIP2 MagicLens EVA-CLIP E5-V VLM2Vec UniME
Classification (10 tasks)
ImageNet-1K 55.8 45.4 10.3 48.0 75.0 40.5 66.5 71.3
N24News 34.7 13.9 36.0 33.7 33.8 31.5 76.4 79.5
HatefulMemes 51.1 47.2 49.6 49.0 49.3 49.3 60.9 64.6
VOC2007 50.7 64.3 52.1 51.6 44.3 76.7 84.0 90.4
SUN397 43.4 39.6 34.5 57.0 62.7 52.3 73.2 75.9
Place365 28.5 20.0 21.5 31.5 38.7 32.0 42.1 45.6
ImageNet-A 25.5 42.6 3.2 8.0 54.8 18.2 39.9 45.5
ImageNet-R 75.6 75.0 39.7 70.9 95.4 56.7 74.6 78.4
ObjectNet 43.4 40.3 20.6 31.6 67.8 34.2 34.3 36.4
Country-211 19.2 14.2 2.5 6.2 38.7 5.9 16.1 18.7
All Classification 42.8 40.3 27.0 38.8 56.0 39.7 56.8 60.6
VQA (10 tasks)
OK-VQA 7.5 2.4 8.7 12.7 9.9 15.1 66.5 68.3
A-OKVQA 3.8 1.5 3.2 2.9 2.8 4.7 54.9 58.7
DocVQA 4.0 4.2 2.6 3.0 7.4 9.1 64.4 67.6
InfographicsVQA 4.6 2.7 2.0 5.9 6.0 8.7 34.8 37.0
ChartQA 1.4 3.0 0.5 0.9 1.5 4.2 33.1 33.4
Visual7W 4.0 1.2 1.3 2.5 2.2 4.5 49.8 51.7
ScienceQA 9.4 7.9 6.8 5.2 14.1 9.6 37.3 40.5
VizWiz 8.2 2.3 4.0 1.7 4.3 8.6 39.9 42.7
GQA 41.3 57.5 9.7 43.5 44.7 34.1 57.3 63.6
TextVQA 7.0 1.0 3.3 4.6 10.8 9.5 65.7 65.2
All VQA 9.1 8.4 4.2 8.3 10.4 10.8 50.4 52.9
Retrieval (12 tasks)
VisDial 30.7 21.5 18.0 24.8 20.4 57.6 75.3 79.7
CIRR 12.6 15.1 9.8 39.1 36.0 41.0 51.3 52.2
VisualNews_t2i 78.9 51.0 48.1 50.7 82.4 43.9 70.7 74.8
VisualNews_i2t 79.6 52.4 13.5 21.1 88.2 46.8 75.2 78.8
MSCOCO_t2i 59.5 58.3 53.7 54.1 65.3 68.6 69.9 74.9
MSCOCO_i2t 57.7 55.0 20.3 40.0 67.2 54.8 67.7 73.8
NIGHTS 60.4 62.9 56.5 58.1 0.2 0.1 63.3 66.2
WebQA 67.5 58.1 55.4 43.0 70.9 33.7 83.6 89.8
FashionIQ 11.4 20.1 9.3 11.2 16.1 11.2 15.2 16.5
Wiki-$S$-NQ 55.0 55.1 28.7 18.7 46.7 61.0 63.4 66.6
OVEN 41.1 56.0 39.5 1.6 1.8 0.5 49.6 55.7
EDIS 81.0 23.6 54.4 62.6 95.6 53.8 73.7 86.2
All Retrieval 53.0 31.6 33.9 35.4 49.2 39.4 63.3 67.9
Visual Grounding (4 tasks)
MSCOCO 33.8 46.4 28.9 22.1 35.8 41.7 77.0 76.5
RefCOCO 56.9 70.8 47.4 22.8 59.9 62.2 85.9 89.3
RefCOCO-matching 61.3 50.8 59.5 35.6 70.0 74.9 83.8 90.6
Visual7W-pointing 55.1 70.1 52.0 23.4 70.2 61.8 83.6 84.1
All Visual Grounding 51.8 59.5 47.0 26.0 58.9 60.2 82.6 85.1
Final Score (36 tasks)
All 39.2 35.0 28.0 27.1 43.7 37.5 63.3 66.6
All IND 37.1 32.3 25.3 31.0 38.1 34.2 64.9 68.4
All OOD 38.7 38.0 25.1 23.7 45.6 33.4 53.9 57.9
C.1. Visualization of the Output Distribution
In Figure 7, we further present additional examples that compare the top-k prediction probabilities of subsequent tokens, illustrating the evolution of model behavior across various training stages of our UniME.
C.2. Schematic illustration of Negative Data
Figure 8 provides additional examples of negative samples encountered during training. Our observations indicate that false negatives are more prevalent in textual data than in visual data, attributable primarily to the presence of synonyms.